Solve cubic root(z+2)^-2/3 = 3 (z+2 is to the negative two thirds power)
Answer by stanbon(75887)

Solve (z+2)^(-2/3) = 3
---
Raise both sides to the -3/2 power to get:
---
z+2 = 3^(-3/2)
z+2 = 1/[sqrt(27)]
z = {1/[3sqrt(3)]} -2
---------------------------
Cheers,
Stan H.